OpenStudy (anonymous):

What inequality represents the sentence? The product of a number and 5 is no more than 8. 5n>8 5n<8

OpenStudy (anonymous):

\[5n \le 8\] 5n<8 5n>8 \[5n \ge 8\]

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

"no more than 8" is like saying "less than or equal to 8"

OpenStudy (anonymous):

\[ 5n \le 8 \]

OpenStudy (anonymous):

thank you :)

OpenStudy (jiteshmeghwal9):

Do according the equation as given product of a number & 5< 8. so, let the unkown no. be 'n' then, the equation will be written as follows 5n<8

OpenStudy (anonymous):

thank you so much :)